# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 2.  Model order reduction problem
In this section, we present a general formulation of the MOR problem. We are concerned with two types of non-autonomous dynamical systems. We focus on nonlinear dynamical systems with external inputs, described by a system of ordinary differential equations (ODE):
@@
where $t\in[t_{0},t_{\infty}]$ denotes time, $\bm{x}(t)\in\mathcal{X}$ denote the differential states, $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, $\bm{u}(t)\in\mathcal{U}$, $\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, and $\bm{f}:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ smooth. We call $\mathcal{X}$ the state space, $\mathcal{U}$ the input space, and Eq. 2 a state space model. The dimension $n_{x}$ is the “model order”. The model is “strictly causal”, because $\bm{x}(t)$ depends only on $\bm{u}(\tau)$, $\tau<t$, and thus $\dot{\bm{u}}(t)$ does not appear in Eq. 2.
We further allow semi-explicit differential-algebraic systems of equations (DAE) with external inputs:
@@
where $\bm{x}(t)\in{\mathcal{X}}\subseteq\mathbb{R}^{n_{x}}$, $\bm{y}(t)\in{\mathcal{Y}}\subseteq\mathbb{R}^{n_{y}}$ are the algebraic variables, and $\bm{u}(t)\in{\mathcal{U}}\subseteq\mathbb{R}^{n_{u}}$ are the external inputs. Let ${\mathcal{X}}$, ${\mathcal{Y}}$ and ${\mathcal{U}}$ be open, and ${\bm{f}}:{\mathcal{X}}\times{\mathcal{Y}}\times{\mathcal{U}}\rightarrow\mathbb{R}^{n_{x}}$ and ${\bm{g}}:{\mathcal{X}}\times{\mathcal{Y}}\times{\mathcal{U}}\rightarrow\mathbb{R}^{n_{y}}$ be smooth. We say that the DAE has “index one” if $\mathrm{D}_{\bm{y}}{\bm{g}}(\hat{\bm{x}},\hat{\bm{y}},\hat{\bm{u}})$ has rank $n_{y}$ for any $(\hat{\bm{x}},\hat{\bm{u}})\in\hat{\mathcal{X}}\times\hat{\mathcal{U}}$ on the solution set of ${\bm{g}}$ [4, 71].
The class of semi-explicit DAEs covers a wide range of models within PSE, where ODEs are often not sufficiently versatile. However, MOR methods are typically developed for ODEs. Fortunately, this is not a problem because an index-one DAE (3) may be regarded as an ODE that lives on a submanifold of $\hat{\mathcal{X}}\times\hat{\mathcal{Y}}\times\hat{\mathcal{U}}$ [72, 15]. Applying definition 1 to Eq. 3b verifies this claim. Moreover, higher-index DAEs can usually be reduced to index-one systems [4]. Consequently, we can adopt ODE methods to index-one DAEs [73, 47]. For conciseness, we will only consider FOMs of the form (2). However, the methods can also be applied to Eq. 3 under the assumption of index-one.
The second reason for introducing both Eqs. 2 and 3 is that some MOR methods reduce a high-order ODE to a low-order DAE, which we generally assume to be index-one. Hence, we will frequently encounter ROMs of the form Eq. 3. Next, we formalize the MOR problem for nonlinear systems with external inputs:
@@
where differential states $\bm{z}(t)\in\mathcal{Z}$, $\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}$, $n_{z}\ll n_{x}$, algebraic variables $\bm{x}_{\mathrm{r}}(t)\in\mathcal{X}$ that approximate the FOM states $\bm{x}(t)$, and $\bm{f}_{\mathrm{r}}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{z}}$, and $\bm{h}:\mathcal{Z}\times\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ are smooth. Let Eq. 4 be such that for any $\bm{x}(t_{0})\in\mathcal{X}$ and any $\bm{u}(t)\in\mathcal{U}$, $t\geq t_{0}$, the prediction error $\left\lVert\bm{x}_{\mathrm{r}}(t)-\bm{x}(t)\right\rVert<\varepsilon$ holds for all $t\geq t_{0}$ and $\varepsilon\geq 0$ sufficiently small.
We call Eq. 4a the “latent dynamics” or “internal dynamics”. A suitable value of $n_{z}$ depends on the FOM, the relevant time scales, and the desired ROM accuracy. For simplicity, we do not impose further requirements on the upper bound $\varepsilon$, e.g., existence of a contraction mapping.
Next, we characterize the subset(s) of $\mathcal{X}$ on which the ROM (4) evolves. To this end, we examine the level set:
$\mathcal{M}\triangleq\{(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}})\in\mathcal{X}\times\mathcal{Z}\times\mathcal{U}:\bm{0}=\bm{h}(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}})\}.$ (5)
According to definition 1, $\mathcal{M}$ constitutes a smooth embedded submanifold of $\mathcal{X}\times\mathcal{Z}\times\mathcal{U}$, where $\dim\mathcal{M}=n_{z}+n_{u}$. In addition, when fixing $\bm{u}_{0}\in\mathcal{U}$, the level set:
$\mathcal{M}_{\mathcal{X}}(\bm{u}_{0})=\{\hat{\bm{x}}\in\mathcal{X}:\bm{0}=\bm{h}(\hat{\bm{x}},\hat{\bm{z}},\bm{u}_{0}),\hat{\bm{z}}\in\mathcal{Z}\}$ (6)
constitutes a smooth embedded $n_{z}$-submanifold of $\mathcal{X}$, where $\mathcal{M}_{\mathcal{X}}(\bm{u}_{0})\times\{\bm{u}_{0}\}$ is a slice of $\mathcal{M}$. To satisfy the requirements in 1, we search for an invariant manifold $\mathcal{M}$ having the property that $\bm{f}(\bm{x}(t),\bm{u}(t))$ is tangent to $\mathcal{M}_{\mathcal{X}}(\bm{u}(t))$ if $\bm{x}(t)\in\mathcal{M}_{\mathcal{X}}(\bm{u}(t))$. Then, Eq. 4 solves the invariance equations (closure condition):
@@
where $\mathrm{D}_{\bm{x}}\bm{h}(\,\cdot\,)^{-1}$ exists due to the index-one assumption and we have used $\mathrm{D}_{\bm{x}}\bm{h}(\,\cdot\,)^{-1}\mathrm{D}_{\bm{u}}\bm{h}(\,\cdot\,)\equiv\mathbf{0}$, since $\bm{f}$ cannot depend on $\dot{\bm{u}}(t)$ due to strict causality.
A fundamental difference to autonomous systems is that changes in $\bm{u}(t)$ can affect $\mathcal{M}_{\mathcal{X}}$. For example, changing $\bm{u}(t)$ may rotate or shift the respective embedded $\mathcal{M}_{\mathcal{X}}$. If $\bm{u}(t)$ is piecewise constant, then $\mathcal{M}_{\mathcal{X}}$ may change its shape abruptly at the input grid points. If $\bm{u}(t)$ is continuous, then $\mathcal{M}_{\mathcal{X}}$ may gradually deform over time. When changing $\bm{u}(t)$ from $\bm{u}_{0}$ to $\bm{u}_{1}$ at time $t_{1}$, the prior $\bm{x}_{\mathrm{r}}^{-}\triangleq\lim_{t\nearrow t_{1}}\bm{x}_{\mathrm{r}}(t)$ may not be in $\mathcal{M}_{\mathcal{X}}(\bm{u}_{1})$. This offset is immediately addressed by a projection of $\bm{x}_{\mathrm{r}}^{-}$ onto $\mathcal{M}_{\mathcal{X}}(\bm{u}_{1})$, inherently defined by Eq. 4b. This projection corresponds to the fast-time-scale response of the system to changes in $\bm{u}(t)$. However, depending on $\bm{h}$, the associated projection can introduce unwanted artifacts in the ROM response (see discussion in Section 3.1.2).
A common simplification of Eq. 4 is to use a one-to-one immersion $\bm{h}^{\dagger}:\mathcal{Z}\rightarrow\mathcal{X}$, so that $\mathcal{M}_{\mathcal{X}}=\{\bm{h}^{\dagger}(\bm{z}):\bm{z}\in\mathcal{Z}\}$ and $\dot{\bm{x}}_{\mathrm{r}}(t)\in\mathrm{col}(\mathrm{D}\bm{h}^{\dagger}(\bm{z}(t)))$ on $\mathcal{M}$. In this case, we have $\mathcal{M}=\mathcal{M}_{\mathcal{X}}\times\mathcal{U}$ (trivial bundle) and Eq. 7 becomes:
@@ 3.  General-purpose reduction methods
We review state-of-the-art reduction methods that are applicable to a wide range of systems, rather than developed for a specific type of process model. Additionally, in Section 3.1.5, we develop an extension of the manifold-Galerkin method (reviewed in Section 3.1.4) to explicitly account for the dependence of the slow manifold on the inputs. For the sake of brevity, we omit equations as far as possible and rather focus on an abstract discussion.
@@ 3.1.  Intrusive general-purpose methods
While POD belongs to the class of linear subspace methods, SPT constructs a nonlinear subspace. Linear subspace methods are typically divided into methods based on singular values decomposition (SVD) and Krylov subspace methods [74]. Herein, we do not discuss Krylov subspace methods [75], which are major reduction techniques for linear systems, but their extension to nonlinear MIMO systems is not straightforward. We refer to [29, 52] for some works towards this goal. An extensive overview of MOR using Krylov subspace methods is given in [18].
@@ 3.1.1.  POD-Galerkin method
The POD, also known as Karhunen-Loéve decomposition, is a method to extract a low-dimensional linear basis capturing the dominant behavior of a system [76]. Usually, the POD basis is determined through SVD of discrete state trajectory samples (“snapshots”) from simulation studies. Define the data matrix $\mathbf{D}\in\mathbb{R}^{n_{x}\times N}$, $N\geq n_{x}$, by stacking $N$ snapshots $\bm{x}^{(k)}$, $k=1,2,...,N$. Compute the mean $\bar{\bm{x}}\in\mathbb{R}^{n_{x}}$ and form the zero-centered matrix $\mathbf{X}=[\bm{x}^{(1)}-\bar{\bm{x}},...,\bm{x}^{(N)}-\bar{\bm{x}}]$. Then, SVD provides:
@@
where $\mathbf{\Sigma}\in\mathbb{R}^{n_{x}\times n_{s}}$ is a non-square diagonal matrix of ordered entries $\sigma_{i}$, which are the “singular values”. The orthogonal matrices $\mathbf{U}\in\mathbb{R}^{n_{x}\times n_{x}}$ and $\mathbf{V}\in\mathbb{R}^{N\times N}$ collect the corresponding left and right singular (column) vectors, respectively. The columns of $\mathbf{U}$ are also called “principal components” of $\mathbf{X}$, since they are ordered by their relevance to represent the data $\mathbf{X}$. As opposed to the eigendecomposition of $\mathbf{X}$, the SVD is well-conditioned and stable [20]. However, SVD is sensitive to data scaling, so improper scaling can result in a poor ROM [76].
The “POD-Galerkin method” consists of two main steps. First, SVD provides the “POD modes” $\mathbf{U}$. Next, we specify the reduced order $n_{z}<n_{x}$ and subdivide $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2}]$, where $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$. Because $\mathbf{U}_{1}$ carries the dominant POD modes, we may truncate the mode $\mathbf{U}_{2}$ and use the approximation:
$\bm{x}^{(k)}-\bar{\bm{x}}\approx\mathbf{U}_{1}\bm{z}^{(k)},\;k=1,2,...,N,$ (10)
where $\bm{z}^{(k)}\in\mathbb{R}^{n_{z}}$ are lower-dimensional coordinates and $\bar{\bm{x}}\in\mathbb{R}^{n_{x}}$ is the mean used in the SVD above. Eq. 10 corresponds to the optimal (“proper”) approximation of the data $\mathbf{X}$ in the 2-norm [77]. However, the optimality of $\mathbf{U}_{1}$ refers only to the data set and not to ROM prediction accuracy [22]. Next, the affine subspace, $\mathcal{M}_{\mathcal{X}}=\{\bar{\bm{x}}+\bm{v}:\bm{v}\in\mathrm{col}(\mathbf{U}_{1})\}$, is defined as the reduced state space. Using Eq. 10, we perform a Galerkin projection of the FOM (2) onto the basis $\mathbf{U}_{1}$ and obtain the ROM [22]:
@@
The POD is data-driven and therefore non-intrusive, but subsequent Galerkin projection is intrusive. Combining data-driven reduced basis and physics-based FOM, the nonlinear ROM (11) is a hybrid model. As the original map $\bm{f}$ is still contained in ROM (11), an additional hyperreduction step may be required to enable real-time applications [41]. In particular, the matrix $\mathbf{U}_{1}$ is usually dense, so a sparse model $\bm{f}$ may lose its sparsity.
Due to CPU costs of SVD, the POD-Galerkin method is limited to problems with a few thousand states [20]. For bigger problems, sequential “greedy” approaches are a cheaper option [78, 34]. These methods minimize the (worst case) $\infty$-norm of the error between reduced basis and snapshots rather than the least squares error or $2$-norm. The reduced basis is then generally non-orthogonal and a projection onto such a basis is a “Petrov-Galerkin projection” [20]. Often, ROMs based on POD and greedy search reach a similar accuracy [79].

### v2

@@ 2. Model order reduction problem
In this section, we present a general formulation of the MOR problem. We are concerned with two types of non-autonomous dynamical systems with inputs. We focus on nonlinear dynamical systems with external inputs, described by a system of ordinary differential equations (ODE):
@@
where $t\in[t_{0},t_{\infty}]$ denotes time, $\bm{x}(t)\in\mathcal{X}$ denote the differential states, $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, $\bm{u}(t)\in\mathcal{U}$ are the system inputs, $\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, and $\bm{f}:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ is smooth. We call $\mathcal{X}$ the state space, $\mathcal{U}$ the input space, and (2) a state space model. Notably, the input trajectory is generally unknown, i.e., $\bm{u}(t)$ are unknown time-variant parameters. The dimension $n_{x}$ is the model order.
We further consider semi-explicit differential-algebraic systems of equations (DAE) with external inputs:
@@
[…]
The class of semi-explicit DAEs covers a wide range of process models, where ODEs are often not sufficiently versatile. However, MOR methods are typically presented in the ODE framework. Fortunately, this is not a problem because an index-one DAE (3) may be regarded as an ODE that lives on a submanifold of $\hat{\mathcal{X}}\times\hat{\mathcal{Y}}\times\hat{\mathcal{U}}$ [239, 203]. Applying Definition 1 to (3b) verifies this point of view. Moreover, higher-index DAEs can usually be reduced to index-one systems [33]. Consequently, we can adopt ODE methods to index-one DAEs, cf. also [170, 114]. Further works that explicitly deal with DAEs are, e.g., in [160, 275, 271, 164, 287]. For conciseness, we will only consider FOMs of the form (2). However, the MOR methods can also be applied to (3) under the assumption of index-one.
The second reason for introducing both (2) and (3) is that some MOR methods reduce a high-order ODE to a low-order DAE, which we generally assume to be index-one. Hence, we will frequently encounter ROMs of the form (3). Next, we formalize the MOR problem for nonlinear systems with external inputs:
@@
where differential states $\bm{z}(t)\in\mathcal{Z}$, $\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}$, $n_{z}\ll n_{x}$, algebraic variables $\bm{x}_{\mathrm{r}}(t)\in\mathcal{X}$ that approximate the FOM states $\bm{x}(t)$, and where $\bm{f}_{\mathrm{r}}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{z}}$ and $\bm{h}:\mathcal{Z}\times\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ are smooth. Let (4) be such that for any $\bm{x}(t_{0})\in\mathcal{X}$, $\bm{u}(t_{0})\in\mathcal{U}$, and $\bm{u}(t){=\hat{\bm{u}}}\in\mathcal{U}$, $t>t_{0}$, the prediction error, $\varepsilon(t)\triangleq\left\lVert\bm{x}_{\mathrm{r}}(t)-\bm{x}(t)\right\rVert$, satisfies $\varepsilon(t)<\varepsilon^{*}$ for all $t>t_{1}$, where $t_{1}\geq t_{0}$ and $\varepsilon^{*}\geq 0$ are sufficiently small.
We call (4a) the latent dynamics or internal dynamics of the ROM. For simplicity, we do not impose further requirements on the upper bound $\varepsilon^{*}$, e.g., the existence of a contraction mapping [10, 140]. In general, a quickly decaying $\varepsilon(t)$ also depends on the ROM initialization strategy, where naïve minimization of $\lVert\bm{x}_{\mathrm{r}}(t_{0})-\bm{x}(t_{0})\rVert_{2}$ may result in a suboptimal $\bm{z}(t_{0})$ [240, 210].
The dimension $n_{z}$ is the reduced model order. A suitable value of $n_{z}$ is fundamentally system- and application-dependent and no universal rule can be given. More specifically, the feasible degree of reduction depends on factors such as i) the type of dynamical system and its attractors (e.g., multiplicity, stability/hyperbolicity), ii) the chosen reduction method, iii) the required accuracy and computational costs of the ROM, and iv) the time-scales/frequencies and subset of state-space relevant to application (e.g., prediction horizon, time steps, bounded region on controllable manifold). For example, some reduction methods provide specific requirements, e.g., on the slow/fast time-scale separation [104, 268], thus limiting the feasible degree of reduction.
Next, we characterize the subset of $\mathcal{X}$ on which the ROM (4) evolves. To this end, we first examine the level set:
[…]
Due to the index-one assumption in Definition 1, $\mathcal{M}$ constitutes a smooth embedded submanifold of $\mathcal{X}\times\mathcal{Z}\times\mathcal{U}$, where $\dim\mathcal{M}=n_{z}+n_{u}$. In addition, when fixing $\hat{\bm{u}}\in\mathcal{U}$, the level set:
[…]
$\mathcal{M}_{\mathcal{X}}(\hat{\bm{u}})\triangleq\{\hat{\bm{x}}\in\mathcal{X}:\bm{0}=\bm{h}(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}}),\;\hat{\bm{z}}\in\mathcal{Z}\}.$ (7)
To satisfy the requirements in Problem 1, we search for an invariant manifold $\mathcal{M}$ having the property that $\bm{f}(\bm{x}(t),\bm{u}(t))$ is tangent to $\mathcal{M}_{\mathcal{X}}(\bm{u}(t))$ if $\bm{x}(t)\in\mathcal{M}_{\mathcal{X}}(\bm{u}(t))$. Then, Eq. (4) solves the invariance equations:
@@
which can be derived by differentiating (4b) with respect to $t$ and inserting (2) and (4a) for $\bm{u}(t)=\hat{\bm{u}}$ fixed. Finally, notice that invariance is closely connected to conservation principles, e.g., the conservation of mass and energy [36, 307].
There exist various local and global approaches for determining invariant (sub)manifolds of nonlinear dynamical systems [151, 281, 130]. Within many of these approaches, the invariance equations (8) play a key role. For example, the parameterization method [40, 110, 109] and extensions [45, 284, 208, 287] have been proposed as a local intrusive method to construct invariant manifolds. If a rigorous solution of the invariance equations is impractical, the invariant manifold may be approximated based on system data (see Section 3.1.4). Finally, projecting the FOM onto an invariant manifold enables a reduced-order representation valid on this subspace.
A fundamental difference to autonomous systems is that changes in $\bm{u}(t)$ can affect $\mathcal{M}_{\mathcal{X}}$. For example, changing $\bm{u}(t)$ may rotate or shift the respective embedded $\mathcal{M}_{\mathcal{X}}$. If $\bm{u}(t)$ is piecewise constant, then $\mathcal{M}_{\mathcal{X}}$ may change its shape abruptly at the input grid points. If $\bm{u}(t)$ is continuous, then $\mathcal{M}_{\mathcal{X}}$ may gradually deform over time. When changing $\bm{u}(t)$ from $\hat{\bm{u}}^{-}$ to $\hat{\bm{u}}^{+}$ at time $t_{1}$, the prior $\bm{x}_{\mathrm{r}}^{-}\triangleq\lim_{t\nearrow t_{1}}\bm{x}_{\mathrm{r}}(t)$ may not be in $\mathcal{M}_{\mathcal{X}}(\hat{\bm{u}}^{+})$. In the ROM, this offset is immediately addressed by a projection of $\bm{x}_{\mathrm{r}}^{-}$ onto $\mathcal{M}_{\mathcal{X}}(\hat{\bm{u}}^{+})$, inherently defined by (4b). Accordingly, the ROM can exhibit a direct feedthrough from the inputs $\bm{u}(t)$ to the reconstructed states $\bm{x}_{\mathrm{r}}(t)$, as indicated by (4b). This projection corresponds to the fast-time-scale response of the FOM to changes in $\bm{u}(t)$. Furthermore, depending on $\bm{h}$, the associated projection after a change in $\bm{u}(t)$ can introduce unwanted artifacts in the ROM response (see Section 3.1.2).
[…]
@@ 3. General-purpose reduction methods
We review state-of-the-art reduction methods that are applicable to a wide range of systems, rather than developed for a specific type of process model. Additionally, in Section 3.1.5, we develop an extension of the manifold-Galerkin method (reviewed in Section 3.1.4) to explicitly account for the dependence of the slow manifold on the inputs. For the sake of brevity, we omit equations as far as possible and rather focus on an informal discussion.
@@ 3.1. Intrusive general-purpose methods
The POD belongs to the class of linear subspace methods, whereas SPT constructs a nonlinear subspace. Linear subspace methods are typically divided into methods based on singular values decomposition (SVD) and Krylov subspace methods [6]. Herein, we do not discuss Krylov subspace methods [12], which are major reduction techniques for linear systems, but their extension to nonlinear MIMO systems is not straightforward. We refer to [226, 9] for some recent works towards this goal. An extensive overview of MOR using Krylov subspace methods is given in [18].
@@ 3.1.1. POD-Galerkin method
The POD, also known as Karhunen-Loéve decomposition, is a method to extract a low-dimensional linear basis capturing the dominant behavior of a system [46]. Usually, the POD basis is determined through SVD of discrete state trajectory samples (snapshots) from simulation studies. Define the data matrix $\mathbf{D}\in\mathbb{R}^{n_{x}\times N}$, $N\geq n_{x}$, by stacking snapshots $\bm{x}(t_{k})$, k=1,2,,Nk=1,2,...,N. Compute the mean $\bar{\bm{x}}\in\mathbb{R}^{n_{x}}$ and form the zero-centered matrix $\mathbf{X}=[\bm{x}(t_{1})-\bar{\bm{x}},...,\bm{x}(t_{N})-\bar{\bm{x}}]$. Then, SVD provides:
@@
where $\mathbf{\Sigma}\in\mathbb{R}^{n_{x}\times N}$ is a non-square diagonal matrix of ordered entries $\sigma_{i}$, which are the singular values. The orthogonal matrices $\mathbf{U}\in\mathbb{R}^{n_{x}\times n_{x}}$ and $\mathbf{V}\in\mathbb{R}^{N\times N}$ collect the corresponding left and right singular vectors, respectively. The columns of $\mathbf{U}$ are also called principal components of $\mathbf{X}$, since they are ordered by their relevance to represent the data $\mathbf{X}$. As opposed to the eigendecomposition of $\mathbf{X}$, the SVD is well-conditioned and stable [5]. However, SVD is sensitive to data scaling, so improper scaling can result in a poor ROM [46].
The POD-Galerkin method consists of two main steps. First, SVD provides the POD modes $\mathbf{U}$. Next, we specify the reduced order $n_{z}<n_{x}$ and subdivide $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2}]$, where $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$. Because $\mathbf{U}_{1}$ carries the dominant POD modes, we may truncate the mode $\mathbf{U}_{2}$ and use the approximation:
$\bm{x}(t_{k})-\bar{\bm{x}}\approx\mathbf{U}_{1}\bm{z}(t_{k}),\;k=1,2,...,N,$ (11)
where $\bm{z}(t_{k})\in\mathbb{R}^{n_{z}}$ are lower-dimensional coordinates and $\bar{\bm{x}}\in\mathbb{R}^{n_{x}}$ is the mean used in the SVD above. Eq. (11) corresponds to the optimal, i.e., proper, approximation of the data $\mathbf{X}$ in the 2-norm [22]. However, the optimality of $\mathbf{U}_{1}$ refers only to the training data and not to ROM prediction accuracy [235]. The reduced state space is defined as the affine subspace, $\mathcal{M}_{\mathcal{X}}=\{\bar{\bm{x}}+\bm{v}:\bm{v}\in\mathrm{col}(\mathbf{U}_{1})\}$. Using (11), we perform a Galerkin projection of the FOM (2) onto the basis $\mathbf{U}_{1}$ and obtain the ROM [235]:
@@
The POD is data-driven and therefore non-intrusive, but subsequent Galerkin projection is intrusive. Combining data-driven reduced basis and physics-based FOM, the nonlinear ROM (12) is a hybrid model. As the original map $\bm{f}$ is still contained in ROM (12), an additional hyperreduction step may be required to enable real-time applications [30]. In particular, the matrix $\mathbf{U}_{1}$ is usually dense, so a sparse model $\bm{f}$ may even lose its sparsity.
For very high-dimensional problems, sequential greedy approaches are computationally cheaper than SVD [29, 117]. These methods minimize the $\infty$-norm of the error between reduced basis and snapshots rather than the least squares error or $2$-norm. The reduced basis is then generally non-orthogonal and a projection onto such a basis is a Petrov-Galerkin projection [5]. Often, ROMs based on POD and greedy search can reach a similar accuracy [250].